OpenStudy (anonymous):

Henry and Irene working together can wash all the windows of their house in 5 h 40 min. Working alone, it takes Henry 8.5 h more than Irene to do the job. How long does it take each person working alone to wash all the windows?

OpenStudy (mertsj):

\[\frac{1}{x}+\frac{1}{x+8.5}=\frac{1}{5\frac{2}{3}}\]

OpenStudy (mertsj):

Solve that. You will get that x = 8.5 and that is how long it takes Irene.

OpenStudy (anonymous):

Which would cause Henry to take 17 Hours correct?

OpenStudy (radar):

Usually with only two workers, you can use "product over sum" x = Irene's work time, and x+8.5 is Henry work time. (x(x+8.5))/(x+x(8.5) = 5 2/3 hrs

OpenStudy (radar):

Yes Henry takes 16.99998 working by himself. He better stick with Irene lol

OpenStudy (radar):

It works out to be a quadratic

OpenStudy (mertsj):

Yep